FINITE TYPE CONDITIONS ON REINHARDT DOMAINS

Siqi Fu Alexander V. Isaev Steven G. Krantz

Abstract. In this paper we prove that, if p is a boundary point of a smoothly bounded pseudoconvex Reinhardt domain in C^{n}, then the variety type at p is identical to the regular type.

Typeset by AMS-TEX